OpenStudy (anonymous):

What percent of 10 is 6?

OpenStudy (anonymous):

6/10% = 0.6% 6/10 = 3/5 = 60% a difference of 4%

OpenStudy (anonymous):

60 % I think because 6/10 is 0.6 = 60%

OpenStudy (dean.shyy):

Put the question into a mathematical equation. So, X% * 10 = 6, and calculate for the X% variable by dividing 6 by 10.